# Semantic Alignment of Multilingual Knowledge Graphs via Contextualized Vector Projections

### v1

@@ I Introduction and Motivation
Knowledge graphs (KGs) are used to represent knowledge bases in a structured way. They use a graph model, i.e., nodes and edges, to represent raw data in a structured format. Multiple applications such as search engines, digital assistants, and semantic web services employ knowledge graphs to store data. Matching corresponding entities across multilingual KGs is challenging due to lexical ambiguity, structural dissimilarity, and linguistic variety. Standard techniques that use syntactic similarity (e.g., Levenshtein, Jaccard) or graph topology do not perform well due to limited overlap of the lexicon or due to the heterogeneous nature of KGs [euzenat2007ontology]. Multilingual KGs such as DBpedia and BabelNet contain cross-lingual data; however, these databases are inconsistent with significant alignment gaps [ehrlinger2016towards]. Manual alignment is not scalable because of its extremely high cost.
We present an embedding-based projection method for aligning multilingual KGs. We use transformer-based pretrained multilingual models, such as SBERT, LaBSE, etc., to generate the embeddings of the entities. The entities are made contextually rich by verbalizing their ontological descriptions. The approach is motivated by three key factors: semantic context (like hierarchy) resolves ambiguity, sentence-level embeddings offer richer meaning than individual tokens, and multilingual LLMs generalize better across languages with different syntax. We further enrich entities with implicit contexts by using HermiT reasoner [horrocks2007hermit] before generating the embeddings. We use cosine similarity and the Hungarian algorithm [kuhn1955hungarian] to find optimal entity mappings, with thresholding to ensure high confidence. This method does not depend on having a specific language as a base language. Also, it does not need a parallel corpus or manual rules to aid its implementation. We have evaluated the proposed method against standard benchmark datasets. Finally, the results have shown that F1 scores are improved relative to lexical, and translation-based baselines.
@@ II Related Work
Traditional methods, such as LogMap [jimenez2011logmap] and AML [faria2013agreementmakerlight], maintain structure, but are inaccurate in ontology matching and have low recall due to translation noise and limited overlap of vocabulary. Methods for representation learning, such as MTransE [chen2017multilingual], which maps embeddings across languages, cannot tackle structural dissimilarities. On the other hand, GNN-based methods like GCN-Align [wang2018cross] capture the local neighborhood but fail on inconsistent subgraphs. Recent works have been following a trend of aligning semantics utilizing entity descriptions (or verbalizations) with multilingual transformers, such as LaBSE [feng2022language], thus breaking language barriers through zero-shot alignment. Embedding-based methods allow for better scalability and better performance on large multilingual knowledge graphs. While recent agent-based Large Language Model (LLM) frameworks offer enhanced reasoning for ambiguous cases [13], vector-projection methods currently provide the necessary scalability and recall for aligning large-scale multilingual datasets.
@@ III Embedding-Based Alignment Architecture
We use multilingual models (mBERT, LaBSE, and XLM-R) [feng2022language] to generate embeddings in the shared multilingual space.
@@ III-A Extended Baseline Comparison
  • Structural Alignment Methods: We used LogMap [jimenez2011logmap] and AML [faria2013agreementmakerlight] as representatives of structure-based alignment systems using ontological reasoning and graph topology.
  • Hybrid Embedding Approaches: We looked over recent methods such as BERT-INT [tang2020bert] that merge contextual embeddings with structural information.
@@ IV Multilingual Semantic Verbalization Strategies
Managing differences in labels across multi-languages has proven to be challenging in multilingual KG alignment [euzenat2007ontology, meilicke2012multifarm]. Since entity names cannot always be directly translated, systems must instead generate rich verbalizations that describe each concept in a language- and context-sensitive manner [sorg2012]. For multilingual models such as LaBSE [feng2022language], XLM-R [conneau2020unsupervised], and mBERT [devlin2019bert], we generate natural-language descriptions of entities (URIs, labels, comments). Short or ambiguous labels (like the generic German term ‘Produkt’) are clarified using the class hierarchy, attributes, and sibling nodes [sun2017cross, chen2017multilingual]. For instance, the DBpedia ontology defines a university as an institution that offers higher education and academic degrees.
Verbalizations mirror shared logical axioms like rdfs:subClassOf and owl:equivalentClass for consistency [horrocks2007hermit]. Consequently, because both are subclasses of dbo:EducationalInstitution, dbo:University and its German counterpart dbr:Universität are defined in this way. Template-based natural language generation can aid this process using the following patterns [Gracia2013cidercl]: Template: A [label] is a [parent class] which is the [attribute description]… The templates use entity metadata for consistent phrasal translations of entities.
When models like XLM-R are fed with contextual enrichment from Wikipedia abstracts, they get more semantic content to work with [sorg2012]. Tokenization methods like SentencePiece keep morphological variations—more specifically, in gendered languages—as is [kudo2018sentencepiece]. Word sense disambiguation methods utilize the context of the entity and the structure of the KG to disambiguate the case where labels are ambiguous [navigli2009word]. It can be made more robust by adding negatives (for example, varying financial and geographical senses of “bank”).
Machine translation is leveraged to produce some synonyms for low-resource languages, which humans then review to maintain meaning fidelity [wu2016google]. To check for overlaps in the language, we check for signing and semantic usefulness using BLEU scoring. The core inputs to the embedding pipeline in Section III are these descriptions.
For the alignment of semantically equivalent entities across multilingual knowledge graphs (KGs), strong and robust similarity modeling is crucial and fundamental [tang2020bert, kumar2025semantic]. We calculate cross-lingual similarity via contextualized vector projections, and then use correspondence methods to turn those similarities into final alignments. We use the multilingual transformers, LaBSE [feng2022language] or XLM-R [conneau2020unsupervised], to encode the verbalized entity explanations generated in Section IV. These models embed each description into a high-dimensional space where similar concepts are placed closer to each other [reimers2020making]. To ensure that cosine similarity is directly proportional to angular proximity, the embeddings are normalized to unit length.
@@
where $\vec{h}_{i}$ and $\vec{h}_{j}$ are normalized contextual embeddings. The scores for all possible pairwise applicant alignments can be captured as a similarity matrix $M\in\mathbb{R}^{p\times q}$ [ngomo2020survey].
Matching Strategies: A naive top-1 correspondence often fails due to semantically similar neighbors, “school” vs “university” [Hertling2019melt]. To enhance precision, we use a thresholded bipartite matching algorithm that has mutual top-k agreement and monotonic score decline [kuhn1955hungarian]. This stops near-synonyms from producing false positives. A post matching-specific action is done to restrict the matching of language ontology particular types or classes (for example, type or parent) [jimenez2011logmap, faria2013agreementmakerlight]. Essentially, these filters ensure that aligned entities share structural characteristics in their respective KGs. The whole pipeline, from verbalized input to similarity cos computation and thresholded correspondence, appears in Figure 4 [qiang2024agent].
@@ V-A Datasets and Experimental Setup
    MTransE [chen2017multilingual]: A translational KG embedding model for alignment.

### v2

@@ I Introduction and Motivation
Knowledge graphs (KGs) are used to represent knowledge bases in a structured way. They use a graph model, i.e., nodes and edges, to represent raw data in a structured format. Multiple applications such as search engines, digital assistants, and semantic web services employ knowledge graphs to store data. Matching corresponding entities across multilingual KGs is challenging due to lexical ambiguity, structural dissimilarity, and linguistic variety. Standard techniques that use syntactic similarity (e.g., Levenshtein, Jaccard) or graph topology do not perform well due to limited overlap of the lexicon or due to the heterogeneous nature of KGs [5]. Multilingual KGs such as DBpedia and BabelNet contain cross-lingual data; however, these databases are inconsistent with significant alignment gaps [4]. Manual alignment is not scalable because of its extremely high cost.
We present an embedding-based projection method for aligning multilingual KGs. We use transformer-based pretrained multilingual models, such as SBERT, LaBSE, etc., to generate the embeddings of the entities. The entities are made contextually rich by verbalizing their ontological descriptions. The approach is motivated by three key factors: semantic context (like hierarchy) resolves ambiguity, sentence-level embeddings offer richer meaning than individual tokens, and multilingual LLMs generalize better across languages with different syntax. We further enrich entities with implicit contexts by using HermiT reasoner [10] before generating the embeddings. We use cosine similarity and the Hungarian algorithm [14] to find optimal entity mappings, with thresholding to ensure high confidence. This method does not depend on having a specific language as a base language. Also, it does not need a parallel corpus or manual rules to aid its implementation. We have evaluated the proposed method against standard benchmark datasets. Finally, the results have shown that F1 scores are improved relative to lexical, and translation-based baselines.
@@ II Related Work
Traditional methods, such as LogMap [11] and AML [6], maintain structure, but are inaccurate in ontology matching and have low recall due to translation noise and limited overlap of vocabulary. Methods for representation learning, such as MTransE [1], which maps embeddings across languages, cannot tackle structural dissimilarities. On the other hand, GNN-based methods like GCN-Align [24] capture the local neighborhood but fail on inconsistent subgraphs. Recent works have been following a trend of aligning semantics utilizing entity descriptions (or verbalizations) with multilingual transformers, such as LaBSE [7], thus breaking language barriers through zero-shot alignment. Embedding-based methods allow for better scalability and better performance on large multilingual knowledge graphs. While recent agent-based Large Language Model (LLM) frameworks offer enhanced reasoning for ambiguous cases [13], vector-projection methods currently provide the necessary scalability and recall for aligning large-scale multilingual datasets.
@@ III Embedding-Based Alignment Architecture
We use multilingual models (mBERT, LaBSE, and XLM-R) [7] to generate embeddings in the shared multilingual space.
@@ III-A Extended Baseline Comparison
  • Structural Alignment Methods: We used LogMap [11] and AML [6] as representatives of structure-based alignment systems using ontological reasoning and graph topology.
  • Hybrid Embedding Approaches: We looked over recent methods such as BERT-INT [23] that merge contextual embeddings with structural information.
@@ IV Multilingual Semantic Verbalization Strategies
Managing differences in labels across multi-languages has proven to be challenging in multilingual KG alignment [5, 16]. Since entity names cannot always be directly translated, systems must instead generate rich verbalizations that describe each concept in a language- and context-sensitive manner [21]. For multilingual models such as LaBSE [7], XLM-R [2], and mBERT [3], we generate natural-language descriptions of entities (URIs, labels, comments). Short or ambiguous labels (like the generic German term ‘Produkt’) are clarified using the class hierarchy, attributes, and sibling nodes [22, 1]. For instance, the DBpedia ontology defines a university as an institution that offers higher education and academic degrees.
Verbalizations mirror shared logical axioms like rdfs:subClassOf and owl:equivalentClass for consistency [10]. Consequently, because both are subclasses of dbo:EducationalInstitution, dbo:University and its German counterpart dbr:Universität are defined in this way. Template-based natural language generation can aid this process using the following patterns [8]: Template: A [label] is a [parent class] which is the [attribute description]… The templates use entity metadata for consistent phrasal translations of entities.
When models like XLM-R are fed with contextual enrichment from Wikipedia abstracts, they get more semantic content to work with [21]. Tokenization methods like SentencePiece keep morphological variations—more specifically, in gendered languages—as is [13]. Word sense disambiguation methods utilize the context of the entity and the structure of the KG to disambiguate the case where labels are ambiguous [17]. It can be made more robust by adding negatives (for example, varying financial and geographical senses of “bank”).
Machine translation is leveraged to produce some synonyms for low-resource languages, which humans then review to maintain meaning fidelity [25]. To check for overlaps in the language, we check for signing and semantic usefulness using BLEU scoring. The core inputs to the embedding pipeline in Section III are these descriptions.
For the alignment of semantically equivalent entities across multilingual knowledge graphs (KGs), strong and robust similarity modeling is crucial and fundamental [23, 15]. We calculate cross-lingual similarity via contextualized vector projections, and then use correspondence methods to turn those similarities into final alignments. We use the multilingual transformers, LaBSE [7] or XLM-R [2], to encode the verbalized entity explanations generated in Section IV. These models embed each description into a high-dimensional space where similar concepts are placed closer to each other [20]. To ensure that cosine similarity is directly proportional to angular proximity, the embeddings are normalized to unit length.
@@
where $\vec{h}_{i}$ and $\vec{h}_{j}$ are normalized contextual embeddings. The scores for all possible pairwise applicant alignments can be captured as a similarity matrix $M\in\mathbb{R}^{p\times q}$ [18].
Matching Strategies: A naive top-1 correspondence often fails due to semantically similar neighbors, “school” vs “university” [9]. To enhance precision, we use a thresholded bipartite matching algorithm that has mutual top-k agreement and monotonic score decline [14]. This stops near-synonyms from producing false positives. A post matching-specific action is done to restrict the matching of language ontology particular types or classes (for example, type or parent) [11, 6]. Essentially, these filters ensure that aligned entities share structural characteristics in their respective KGs. The whole pipeline, from verbalized input to similarity cos computation and thresholded correspondence, appears in Figure 4 [19].
@@ V-A Datasets and Experimental Setup
    MTransE [1]: A translational KG embedding model for alignment.